OpenStudy (anonymous):

The pressure of a gas is reduced from 1200.0 mm Hg to 850.0 mm Hg as the volume of its container is increased by moving a piston from 85.0 mL to 350.0 mL. What would the final temperature be if the original temperature was 90.0 °C?

OpenStudy (kendricklamar2014):

Given: P1 = 1200.0 mmHg P2 = 850 mmHg V1 = 85.0 mL V2 = 350.0 mL T1 = 90.0 deg C T2 = ? Convert the temperature from Celcius to Kelvins. K = temperature in Celcius + 273 K = 90.0 deg C + 273 K = Combined Gas Law: P1 x V1 / T1 = P2 x V2 / T2 Rearrange the formula so T2 = P2 x V2 x T1 / P1 x V1 T2 = (850.0 mmHg x 350.0 mL x 90.0 deg C) / (1200.0 mmHg x 85.0 mL)

OpenStudy (anonymous):

omg you're awesome thank u my teacher confuses me

OpenStudy (kendricklamar2014):

No Problem